OpenStudy (anonymous):

Prove the following trigonometric identity: (1-2sin^2x)/(cosx+sinx)=cosx-sinx Thanks for any help.

hartnn (hartnn):

do u know \(\large \sin^2x+\cos^2x=1\) ???

OpenStudy (anonymous):

\[\frac{ 1-2\sin ^{2}x }{ \cos x+\sin x }=\cos x-\sin x\]

OpenStudy (anonymous):

Yes I do

hartnn (hartnn):

then substitute that 1 in the numerator by sin^2 x + cos^2 x what do u get ?

OpenStudy (anonymous):

sin^2x+cos^2x-2sin^2x?

hartnn (hartnn):

yes, what does that simplify to ?

OpenStudy (anonymous):

-sin^2x+cos^2x

hartnn (hartnn):

correct! now thats in \(\large a^2-b^2 \) form , right > can you factorize ?

hartnn (hartnn):

like a^2-b^2 = (a+b)(a-b)

OpenStudy (anonymous):

So (sinx+cosx)(sinx-cosx)?

hartnn (hartnn):

wouldn't the 2nd term be cos x - sinx ? because we have cos^2 x - sin^2 x right ?

OpenStudy (anonymous):

What exactly do you mean by 2nd term?

hartnn (hartnn):

(sinx+cosx)(sinx-cosx) the sin x - cos x term should be cos x - sin x

hartnn (hartnn):

-sin^2x+cos^2x = cos^2 x- sin^2 x = (cos x + sin x)(cos x - sin x) got this ?

OpenStudy (anonymous):

Yes i follow it

hartnn (hartnn):

and what gets cancelled ?

hartnn (hartnn):

from numerator and denominator ?

OpenStudy (anonymous):

cosx+sinx?

hartnn (hartnn):

yes! and what remains ?

OpenStudy (anonymous):

cosx-sinx Thank you so much for your help!

hartnn (hartnn):

welcome ^_^